Determine the mass of 20.0 cubic feet of aluminum whose density is

2.70 g/cm", expressed in kilograms. (b) Diamonds are measured in carats,

and 1 carat = 0.200g. The density of diamond is 3.51 g/cm°. What is the

volume of a 5.0-carat diamond, expressed in liters?


1
Expert's answer
2021-05-10T01:21:27-0400

a. 1 cubic foot = 28.3168466 liters


20 cubic feet = 20*28.3168466 = 566.336932 liters = 566336.932 cm3


m = 2.7g/cm3*566336.932cm3= 1529109.72 g = 1529.11 kg


b mass of diamond =5×0.200g = 1g

Density= mass/Volume

Volume = mass/ density

V= 1g/3.51g/cm³ = 0.28 cm³

Volume in liters = 0.28mL